OpenStudy (anonymous):

(Trigonometry) How to convert 2x - 3y = 5 to polar form?

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

Do you know Trig?

OpenStudy (uri):

You're in which grade?

OpenStudy (anonymous):

Going to be a Senior.

OpenStudy (anonymous):

12th

OpenStudy (uri):

Oh which group? I mean Commerce?

OpenStudy (anonymous):

Sorry, I'm in High School.

OpenStudy (uri):

Oh okay cool,I'm just 12 so no idea about it. :)

OpenStudy (anonymous):

Man... Nobody can answer this...

OpenStudy (uri):

@jim_thompson5910 Can :D

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

hint: x = r*cos(theta) y = r*sin(theta) make sure you solve for r once you've made all the substitutions

OpenStudy (anonymous):

I got as far as 5/r=2cos(theta)-3sin(theta), but I think that it is wrong...

jimthompson5910 (jim_thompson5910):

2x - 3y = 5 2*r*cos(theta) - 3*r*sin(theta) = 5 r*(2*cos(theta) - 3*sin(theta)) = 5 r = 5/(2*cos(theta) - 3*sin(theta)) so you were (kinda) close

jimthompson5910 (jim_thompson5910):

not sure how you got 5/r though

OpenStudy (anonymous):

For this one, would it be r^2=sqrt(10)/2 ?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

x2 + y2 = 10 My bad... That is the starting equation...

OpenStudy (anonymous):

Oh, but it would be jut r=sqrt(10)/2

OpenStudy (anonymous):

Right?

jimthompson5910 (jim_thompson5910):

oh this is a circled centered at (0,0) with radius sqrt(10)

jimthompson5910 (jim_thompson5910):

so it's just r = sqrt(10) since r is the radius

OpenStudy (anonymous):

Oh! I see. I think that I am over-complicating this...

jimthompson5910 (jim_thompson5910):

possibly, but that's what polar form is essentially: each point in the coordinate space is defined by the angle and how far it is from the center think of it like you're on a navy battleship and you are looking at the radar

jimthompson5910 (jim_thompson5910):

and you can think of the angle as the direction you need to look in to see the point

OpenStudy (anonymous):

I think I understand. So for x^2 - y^2 = 4 (You may leave your answer for this one as r2) It is r^2=sqrt(-2), and that is it, right?

jimthompson5910 (jim_thompson5910):

hmm no that's a bit off

OpenStudy (anonymous):

*sigh* I am bad at these...

jimthompson5910 (jim_thompson5910):

x^2 - y^2 = 4 (r*cos(theta))^2 - (r*sin(theta))^2 = 4 r^2*cos^2(theta) - r^2*sin^2(theta) = 4 r^2(cos^2(theta) - sin^2(theta)) = 4 r^2*cos(2theta) = 4 r^2 = 4/cos(2theta) r = sqrt( 4/cos(2theta) )

jimthompson5910 (jim_thompson5910):

its ok, just keep practicing

OpenStudy (anonymous):

All right, well thanks bunches. You gave me a couple examples to go off of :)

jimthompson5910 (jim_thompson5910):

you're welcome